What is the equation in slope-intercept form for m=1/2 y-intercept (4, 3)?
Answer by Edwin McCravy(20055)

What is the equation in slope-intercept form for m=1/2, y-intercept (4, 3)?
Sorry. This problem is botched up because (4, 3) is not a point 
on the y-axis and thus cannot be a y-intercept.

If the problem were either this:

What is the equation in slope-intercept form for m=1/2 y-intercept (0, 3)?

or this:

What is the equation in slope-intercept form for m=1/2 containing
the point (4, 3)?

it would be answerable.  Please repost.

Edwin